40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16272     Entry time: Fri Aug 6 17:10:19 2021
Author: Paco 
Type: Update 
Category: IMC 
Subject: MC rollercoaster 

[anchal, yehonatan, paco]

For whatever reason (i.e. we don't really know) the MC unlocked into a weird state at ~ 10:40 AM today. We first tried to find a likely cause as we saw it couldn't recover itself after ~ 40 min... so we decided to try a few things. First we verified that no suspensions were acting weird by looking at the OSEMs on MC1, MC2, and MC3. After validating that the sensors were acting normally, we moved on to the WFS. The WFS loops were disabled the moment the IMC unlocked, as they should. We then proceeded to the last resort of tweaking the MC alignment a bit, first with MC2 and then MC1 and MC3 in that order to see if we could help the MC catch its lock. This didn't help much initially and we paused at about noon.

At about 5 pm, we resumed since the IMC had remained locked to some higher order mode (TEM-01 by the looks of it). While looking at C1:IOO-MC_TRANS_SUMFILT_OUT on ndscope, we kept on shifting the MC2 Yaw alignment slider (steps = +-0.01 counts) slowly to help the right mode "hop". Once the right mode caught on, the WFS loops triggered and the IMC was restored. The transmission during this last stage is shown in Attachment #1.

Attachment 1: MC2_trans_sum_2021-08-06_17-18-54.png  96 kB  Uploaded Fri Aug 6 18:20:35 2021  | Hide | Hide all
MC2_trans_sum_2021-08-06_17-18-54.png
ELOG V3.1.3-